OpenStudy (anonymous):

Differential Equations: Solve the equation: (1+x^2) dy + (1 + y^2) dx = 0 ; y(1) = 0

OpenStudy (dan815):

separable

OpenStudy (dan815):

use a trig sub

OpenStudy (anonymous):

Question: Is it not an exact equation?

OpenStudy (dan815):

actuallly lets check

OpenStudy (dan815):

dm/dy=dn/dx if its exat

OpenStudy (dan815):

nope its not exact

OpenStudy (dan815):

in general an equal is exact if the equation can be written in the form Grad F

OpenStudy (anonymous):

My = 0 and Nx = 0, right?

OpenStudy (dan815):

no its 2y and 2x

OpenStudy (anonymous):

if M = (1 + x^2) and N = (1 + y^2)

OpenStudy (dan815):

no switch it

OpenStudy (dan815):

notice they wrote the order dy dx

OpenStudy (dan815):

ally its dx then dy

OpenStudy (dan815):

usually*

OpenStudy (dan815):